### v1

# Search for joint multimessenger signals from potential Galactic PeVatrons with HAWC and IceCube

Abstract
Galactic PeVatrons are sources that can accelerate cosmic rays to PeV energies. The high-energy cosmic rays are expected to interact with the surrounding ambient material or radiation, resulting in the production of gamma rays and neutrinos. To optimize for the detection of such associated production of gamma rays and neutrinos for a given source morphology and spectrum, a multi-messenger analysis that combines gamma rays and neutrinos is required. In this study, we use the Multi-Mission Maximum Likelihood framework (3ML) with IceCube Maximum Likelihood Analysis software (i3mla) and HAWC Accelerated Likelihood (HAL) to search for a correlation between 22 known gamma-ray sources from the third HAWC gamma-ray catalog and 14 years of IceCube track-like data. No significant neutrino emission from the direction of the HAWC sources was found. We report the best-fit gamma-ray model and 90% CL neutrino flux limit from the 22 sources. From the neutrino flux limit, we conclude that the gamma-ray emission from five of the sources can not be produced purely from hadronic interactions. We report the limit for the fraction of gamma rays produced by hadronic interactions for these five sources.
1 Introduction
Since the discovery of cosmic rays (CRs) more than 100 years ago (Hess, 1912), but where these high-energy particles come from has not yet been well understood. The highest energy charged particles, protons and nuclei, are thought to be extraterrestrial in origin, and their spectrum can be parametrized using a power law with an index of 2.7 up to PeV (or the “knee” in the spectrum), above which the spectrum steepens (Abbasi et al., 2018). The CRs below the knee are thought to originate from sources within the Milky Way galaxy, hinting at the presence of PeV hadron accelerators within the galaxy known as Galactic PeVatrons (Gabici & Aharonian, 2007). Galactic PeVatrons are expected to accelerate particles up to PeV energies through Fermi acceleration and these accelerated particles are expected to be confined within the Milky Way galaxy (Abbasi et al., 2018).
Supernova remnants (SNRs) are a potential class of Galactic PeVatrons. The particles are accelerated by diffusive shock acceleration in the expanding shock. Only 10% of the energy of SNRs is needed to explain the observed CR energy density in the Galaxy (Blandford & Ostriker, 1978). For the most energetic galactic CRs, it is assumed that only young SNRs can accelerate CRs beyond 100 TeV (Bell et al., 2013). The accelerated CRs can escape the SNR and diffuse into the surrounding interstellar medium. They can interact with the surrounding material like molecular clouds and produce charged and neutral pions. Therefore, young SNRs located near star clusters or molecular clouds are thought to be bright in gamma rays due to hadronic interactions. High-energy gamma rays and neutrinos can be efficiently produced by the decay of charged and neutral pions. Other potential source classes of Galactic PeVatrons include young massive star clusters, YMCs (Bykov et al., 2020) and pulsar wind nebulae, PWNe (Di Palma et al., 2017).
TeV gamma rays can also be produced by leptonic interactions such as inverse Compton scattering. However, inverse Compton scattering is strongly suppressed at energies above TeV (Hinton & Hofmann, 2009) due to the Klein-Nishina effect (Klein & Nishina, 1929). On the other hand, the gamma rays produced by neutral pion decays are not suppressed at O(10 TeV), making TeV gamma-ray emitters excellent candidates for Galactic PeVatrons. Moreover, high-energy neutrinos are also produced in hadronic interactions via charged pion decays and their detection is evidence of hadron acceleration. Because of this relationship, a joint search using TeV gamma-ray and neutrino data from known TeV gamma-ray emitters can test whether a source is a Galactic PeVatron.
Several potential Galactic PeVatrons have been investigated by the community. The HAWC collaboration reported that the TeV gamma-ray spectrum of HAWC J1825-134 extends up to 200 TeV without any cut-off, much higher than the Klein-Nishina suppression energy(Albert et al., 2021). This may indicate hadronic acceleration. The HAWC collaboration also reported HAWC J2227+610 as a potential PeVatron candidate from TeV gamma-ray observations and concluded the proton cutoff energy is at least 800 TeV (Albert et al., 2020a). The HAWC collaboration studied the ultra-high-energy spectrum off MGRO J1908+06 and their model allows a hadronic component at the highest energy (Albert et al., 2022). The HESS Collaboration reported evidence of a PeVatron at the galactic center based on gamma-ray observations up to 40 TeV (Abramowski et al., 2016).
The IceCube Neutrino Observatory has previously searched for neutrino emission from galactic sources, both individual sources and classes of sources. Abbasi et al. (2023a) searched for neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using a fixed spatial extension and a power law spectrum. They found no significant emission, neither from single sources nor from stacked catalogs of SNRs and PWNe. Recently, IceCube has detected neutrinos from the galactic plane (Abbasi et al., 2023b). The observation was made using a neutrino detection channel, cascades, which has relatively large angular uncertainty. Therefore, IceCube could not distinguish between diffuse emission in the galactic plane and a collection of individual sources. The IceCube collaboration also searched the entire galactic plane and regions near TeV gamma-ray sources using an extended source hypothesis and a power law spectrum, finding no significant neutrino emission (Abbasi et al., 2023c). However, these searches have not directly incorporated gamma-ray data. The simultaneous use of both data sets incorporates more physically motivated spectra and provides a direct constraint on the hadronic ratio, i.e. the fraction of gamma rays originating from hadronic interactions.
If the gamma rays and neutrinos originate from hadronic interactions, the spectra of the gamma rays and neutrinos have a fixed relationship. In the case of Galactic PeVatrons, the proton-proton (pp) interaction is expected to be dominant as compared to the photo-hadronic interactions (Ahlers & Murase, 2014). In the pp interactions, protons interact with nearby material and produce charged and neutral pions. The ratio between the production of charged and neutral pions is about 2:1. Charged pions decay further into three neutrinos and other products, with each resulting neutrino having about a quarter of the energy of the charged pion. The neutral pions decay into two gamma rays, each with half of the energy of the neutral pion, on average. Therefore, we expect that for charged and neutral pion decay, the energy of the neutrino produced will be half the energy of the gamma-ray produced on average. On Earth, the observed flux is
where is the neutrino differential particle flux with the neutrino flavor , is the distance to the source, is the interaction length of the gamma ray and is the gamma-ray differential particle flux (Ahlers & Murase, 2014). Since we only consider galactic sources, we assume that the absorption of gamma-rays is negligible () and the flux from each neutrino flavor is the same due to oscillations. The relationship between the gamma-ray flux () and the muon neutrino flux () is thus
Therefore, if the gamma rays from a galactic source originate from hadronic interactions, the neutrino counterpart should exhibit the same morphology and have the same spectral shape shifted to lower energies. This relationship is used in this work to model and simultaneously fit the gamma-ray and neutrino data.
2 Instruments and Datasets
2.1 HAWC Gamma-ray Observatory
The High Altitude Water Cherenkov (HAWC) Observatory is located near Sierra Negra, Mexico, at a latitude of and an altitude of 4100 m. It has been in continuous operation since November 2014 (Abeysekara et al., 2023). The observatory consists of 300 Water Cherenkov Detectors (WCDs), covering an area of 22000 . Each WCD consists of a steel tank 7.3 m in diameter and 5.4 m high, lined with a plastic bladder and filled with purified water. At the bottom of the water tank, three 8-inch photomultiplier tubes (PMTs) are arranged in an equilateral triangle with a side length of 3.2m, and with one 10-inch PMT in the center (Abeysekara et al., 2023).
When high-energy gamma rays interact with the Earth’s atmosphere, they create a cascade of particles called an extensive air shower (EAS). When these particles reach the HAWC detector and pass through the water at a speed greater than the speed of light in water, Cherenkov light is induced and collected by the PMTs in the WCDs. HAWC uses the observed Cherenkov emission to reconstruct the properties of the primary gamma rays.
HAWC divides the detected shower events into a 2D array of analysis bins based on the fraction of triggered PMTs during a shower event and the reconstructed energy of the event (Abeysekara et al., 2019a). The events within an analysis bin are expected to have similar characteristics, including angular resolution, primary gamma-ray energy, and gamma-to-hadron likelihood (signal-to-noise ratio), which allows HAWC to perform a binned likelihood analysis.
In this study, we use 2141 days of HAWC data using the Pass 5 reconstruction (Yun-Carcamo et al., 2023). The event-related energy is estimated using the “ground parameter” estimator described in Abeysekara et al. (2019a). This energy estimator uses the charge at a fixed optimal distance from the shower axis to estimate the energy. The dataset covers the period from June 15, 2015 to October 21, 2021 with a livetime of more than 90%.
2.2 The IceCube Neutrino Observatory
The IceCube Neutrino Observatory is a neutrino detector located at the geographic South Pole (Aartsen et al., 2017). The primary IceCube detector comprises a cubic kilometer array of 5160 digital optical modules (DOMs). These DOMs are arranged along 86 readout and support cables, called strings, embedded in the glacial ice at depths between 1.45 km and 2.45 km. Each DOM consists of a 10-inch PMT and its readout electronics and is housed in a pressure-resistant sphere (Abbasi et al., 2009). DOMs are designed to detect the Cherenkov light produced by charged particles moving through the ice (Abbasi et al., 2010). These particles originate from the charged-current (CC) interactions or neutral-current (NC) interactions of all flavors of neutrinos. IceCube events are typically divided into two categories: track-like events produced by CC interactions of muon neutrinos and cascade-like events from all other CC interactions and NC interactions. The DOMs record the Cherenkov light, allowing events to be classified and event properties, such as the direction of the incident neutrinos and the deposited energy, to be reconstructed (Aartsen et al., 2014).
The dataset used in this analysis includes 14 years of IceCube track-like data from April 6, 2008 to May 23, 2022. Track-like data provide a better angular resolution ( angular uncertainty for a typical TeV neutrino) compared to cascade-like events. This makes track-like events more suited to search for sources with an extension below a few degrees. The data from April 6, 2008 to May 13, 2011 were taken with the partially constructed detector with 40, 59, and 79 strings (IC40, IC59, and IC79), with each configuration spanning approximately one year (Aartsen et al., 2014). The remaining data were recorded with the full detector configuration with 86 strings (IC86).
3 Analysis method
3.1 Source selection
HAWC detected 65 sources with a significance above in the third HAWC catalog (3HWC) (Albert et al., 2020b) using 1523 days of data. The 3HWC provides the location and the gamma-ray flux at a pivot energy of 7 TeV for each source in the catalog. The pivot energy is the energy of normalization of the spectral function, which is specified in terms of . If the gamma-ray emission from these sources comes from hadronic interactions, the neutrino flux will be described by the same spectral function with half the pivot energy (3.5 TeV), according to Equation 2. We calculate the IceCube 90% confidence level (CL) sensitivity for a point source with an unbroken power law spectrum. The sensitivity is defined as the median expected 90% CL upper limit. The sensitivity can be determined using background trials and signal trials. Background trials are done by performing the analysis using simulated data obtained by scrambling real data in right ascension (RA). Scrambling data as background does not suffer from systematic uncertainties and therefore is a robust way to generate background trials. Signal trials are generated by scrambling data and adding signal using simulation data. The calculation of sensitivity is based on a maximum likelihood method as used in the IceCube 10-year search for neutrino sources(Aartsen et al., 2020a). In this work, we use a new likelihood analysis software component called i3mla (Abbasi et al., 2021) with the multi-mission maximum likelihood framework (3ML) (Vianello et al., 2017). Details of i3mla are described in section 3.4. The likelihood consists of a spatial term and an energy term, as explained in more detail in Appendix A.
We select sources whose predicted neutrino flux from hadronic interactions is larger than the sensitivity of the IceCube point source search. The hard spectrum is an optimistic choice made to avoid missing any potential sources in our source selection step. The final sensitivity of the joint search will change depending on the best-fit gamma-ray spectrum and spatial extension. The sensitivity of IceCube and the predicted neutrino flux from 3HWC sources are shown in Figure 1. We exclude two extra-galactic sources, Markarian 421 and Markarian 501, as this study focuses on galactic sources. We also exclude four sources classified as secondary sources in 3HWC, as these are likely to be statistical fluctuations near bright sources. Our final source list comprises 22 sources from 3HWC. Many of the sources are still unidentified but some have potential associations with pulsar wind nebulae, supernova remnants, and star-forming regions.
3.2 Model selection and fit using gamma-ray data
To obtain a spatial and spectral model for the joint analysis of the sources, we first fit each gamma-ray source with an updated HAWC data set. We use 2141 days of HAWC ground parameter energy estimator maps with the Pass 5 reconstruction for our model selection procedure. The dataset contains more data and a better energy estimator compared to the dataset used in 3HWC. The new reconstruction provides a better angular reconstruction and gamma/hadron separation. We use the 3ML framework (Vianello et al., 2017) with the HAWC Accelerated Likelihood (HAL) plugin (Abeysekara et al., 2021) for our model selection and model fitting. HAL is a Python-based, maximum likelihood software module designed to perform HAWC analysis using a binned likelihood. In conjunction with 3ML, it can perform both stand-alone HAWC analyses and joint analyses with other experiments. We select the model that best describes the gamma-ray emission based on the log-likelihood (LLH). We use the position in 3HWC as a starting point for our initial fit. First, we use a simple power law and a point source as spectral and spatial hypotheses. The simple power law is parameterized as , where is the flux normalization and is the spectral index. The free parameters in this model are the Right Ascension (RA), the Declination (Dec), the spectral index, and the gamma-ray flux. Second, we test the extended source hypothesis, where we fix the RA and Dec of the source to the best-fit values and fit for a Gaussian extension along with the spectral index and flux. We accept the extended source hypothesis if twice the difference between the log-likelihood () of the two models (point source hypothesis and extended source hypothesis) is greater than 16, corresponding to according to Wilks’ theorem (Wilks, 1938). Finally, we test the spectrum curvature hypothesis, where we fixed the RA, Dec, and extension (if the source is extended) and we use a log-parabola spectrum to fit the data again. The log-parabola spectrum is parameterized by
The free parameters in this model are , and the gamma-ray flux at 7 TeV. We accept the spectral curvature hypothesis if the between the spectral curvature hypothesis and the power law hypothesis is greater than 16. Based on the results of these fits, we obtain the final model that contains information about the morphology and spectrum of the source.
3.3 IceCube likelihood with i3mla
Combining IceCube neutrino data with the HAWC gamma-ray data is a non-trivial task because of the complexity of the instrument response functions and the likelihood calculation. We use a Python-based maximum likelihood module called i3mla (Abbasi et al., 2021) for IceCube likelihood calculation. Similar to HAL, i3mla is specifically designed to be compatible with the 3ML framework. i3mla uses instrument response functions (IRFs) to encapsulate the detector properties and performances of the detector determined from Monte Carlo simulations. The IRFs then can be used to calculate the energy likelihood of IceCube. This differs from other IceCube analysis software that pre-computes the energy likelihood for an assumed spectrum, e.g. a simple power law, and uses all the Monte Carlo data for the energy likelihood calculation. The new tabulated IRFs method speeds up the likelihood computation process when fitting for an arbitrary spectrum, making it possible to fit any spectral parameters during the model fitting in i3mla.
In this analysis, we use spatial and energy likelihood terms to calculate the likelihood of the model. The signal hypothesis assumes some neutrino emission from the source of interest. The background hypothesis assumes all non-neutrino events and neutrino events originate from atmospheric muons, atmospheric neutrinos, and astrophysical neutrinos that are not associated with the source. We model the IceCube point spread function (PSF) as a 2D Gaussian and the angular error of the 2D Gaussian is estimated during the reconstruction. The signal spatial likelihood is the PSF convolved with the morphology of the source and the signal energy likelihood is calculated using the IRFs and the model spectrum. The background spatial likelihood is estimated from the data by binning the data in reconstructed declination and creating a spline using the resulting histogram. The background spatial likelihood is a function of declination only. The background energy likelihood is estimated from the data by weighting the simulation using the spectral hypothesis and binning it in reconstructed declination and energy. The method is conceptually identical to the previous IceCube 10-year time-integrated point source search in Aartsen et al. (2020a) and the extended source search in Abbasi et al. (2023a, c). A detailed description of the likelihood can be found in Appendix A.
3.4 Joint fit with Gamma rays and Neutrinos
After obtaining a spatial and spectral model for each source, we perform a joint fit with HAWC gamma-ray data and IceCube neutrino data. We keep the location and extension of the model fixed. We add a neutrino component to the model that coincides with the gamma-ray source. We connect the neutrino and gamma-ray emission spectral parameters according to the relation in Equation 2 and allow them to float during the fit. We let the flux normalization of the neutrino and gamma-ray components float independently. Table 1 shows the properties of the model parameters.
| Model parameters | Gamma-ray spectrum | Neutrino spectrum | Relation |
| Right ascension(RA) | Fix | Fix | Equal |
| Declination(dec) | Fix | Fix | Equal |
| Extension | Fix | Fix | Equal |
| Float | Float | Equal | |
| Float | Float | Equal | |
| Pivot energy | Fix | Fix | Neutrino pivot energy = Gamma-ray pivot energy / 2 |
| Flux normalization | Float | Float | Independent |
To calculate the p-value, we perform a large number of background trials by scrambling the IceCube data in RA and building the background test statistic (TS) distribution (Aartsen et al., 2020b). This is done to avoid any dependence on the simulation when testing for a neutrino source. The calculation of the IceCube component of the likelihood is fast. However, using raw HAWC binned maps for computing the HAWC component of the likelihood is prohibitively computationally expensive for such a large number of trials. Therefore, the HAWC data are fixed during the background scrambling process so that we can pre-compute the HAWC component of the likelihood. We extract the HAWC likelihood around the best-fit spectral parameters and save it into a table. Since the location and extension are fixed during the fit, we do not extract the likelihood around the spatial parameters. For spectral parameters, we evaluate the HAWC likelihood within of the best-fit spectral shape parameters while maximizing it over the flux normalization.
Since the gamma-ray and neutrino fluxes are floated independently, maximizing the likelihood over the gamma-ray flux normalization leads to a maximum likelihood for HAWC at the spectral shape parameters in the joint fit. For a log-parabola spectrum, there are two free parameters for the spectral shape ( and ), and we use a 5050 grid around the best-fit parameters for the likelihood evaluation. For a power-law spectrum, there is only one free parameter for the spectral shape () and we evaluate 300 points around the best-fit parameters. We have developed a new custom plugin for 3ML to read the likelihood table and construct a spline. The 3ML plugin returns the spline value for the likelihood evaluation. This method speeds up the likelihood calculation because the program does not have to calculate the HAWC likelihood for each pixel and each bin during the fitting. Figure 2 shows an example of the likelihood.
The background TS distribution is calculated by generating 50,000 background-only trials. For each background trial, the IceCube data are scrambled in RA while keeping the HAWC component of the likelihood unchanged, and the joint analysis is performed. To determine the sensitivity to the neutrino flux of the joint search, we inject different simulated values of the neutrino flux at the location of the best-fit gamma-ray source with its best-fit extension and spectral shape. The 90% CL sensitivity is the neutrino flux required for the TS to exceed the median of the background-only TS distribution 90% of the time. The discovery potential is the neutrino flux required for the TS to exceed the fluctuation of the background-only TS distribution 50% of the time. Figure 3 shows the ratio between the sensitivity and the predicted neutrino flux assuming pp interactions, as well as the ratio between the discovery potential and the predicted neutrino flux assuming pp interactions.
We derive the pre-trial p-value for each source by using the background TS distribution of each source. The pre-trial p-value is the percentage of background trials having a TS higher than the TS of the actual search. To correct for the look-elsewhere effect, we create a pre-trial p-value distribution by searching for neutrinos from each source using the same background data and picking the smallest p-value from the 22 sources in each background trial. The post-trial p-value is the fraction of p-values in the aforementioned distribution that are smaller than the smallest pre-trial p-value in the data. With this method, the correlations between different sources can be correctly taken into account.
In addition to searching for individual sources, we perform a binomial test to determine whether a subset of the sources has TS values inconsistent with the background (Abbasi et al., 2022a). The test statistic of the binomial test, , is the minimum probability, over all that at least p-values are at or below the kth smallest p-value observed, :
To account for the look-elsewhere effect, we create a distribution of by performing the binomial test on scrambled data. We compute the post-trial p-value of the binomial test using the value for data. To calculate the overall post-trial p-value accounting for the trials from doing two searches, we create another p-value distribution by choosing, for each background trial, the smaller of the post-trial p-values from the individual source search and the binomial test. This method takes into account the correlation between the individual source search and the binomial test. We then compare the smaller of the two p-values from data with this distribution to obtain the overall p-value.
4 Results
4.1 HAWC best-fit result
First, we present the best-fit results for the gamma-ray emission from the HAWC Pass 5 data alone. Table 2 summarizes the best-fit location, extension, and spectral parameters of each source in our list. In addition, we report the statistical and systematic uncertainty of the spectral fit results. Systematic uncertainties of the HAWC detector include PMT efficiency, charge uncertainty, etc., and are described in Abeysekara et al. (2019b). To evaluate their effect, we perform the model selection using samples simulated with different settings and compare the results with the baseline simulation. The errors of each type of systematic uncertainty are added in quadrature to obtain the total systematic uncertainty, which is given in the table 2.
| Name | RA [∘] | Dec [∘] | Ext [∘] | [TeV-1cm-2s-1] | Nearest TeVCat source | ||
| 3HWC J0534+220 | 83.64 | 22.01 | PS | 2.82 | 0.12 | Crab | |
| 3HWC J0634+067 | 98.59 | 6.66 | 0.49 | 2.57 | 0 | HAWC J0635+070 | |
| 3HWC J1757240 | 269.32 | 23.86 | PS | 2.58 | 0 | HESS J1800240B | |
| 3HWC J1809190 | 272.42 | 19.34 | 0.23 | 2.09 | 0.23 | HESS J1809193 | |
| 3HWC J1813125 | 273.4 | 12.7 | 0.27 | 2.6 | 0 | HESS J1813-126 | |
| 3HWC J1813174 | 273.4 | 17.74 | 0.26 | 1.91 | 0.35 | 2HWC J1814173 | |
| 3HWC J1825134 | 276.49 | 13.63 | 0.46 | 2.45 | 0.13 | 2HWC J1825134 | |
| 3HWC J1831095 | 277.85 | 9.85 | 1.14 | 2.59 | 0.14 | HESS J1831098 | |
| 3HWC J1837066 | 279.38 | 6.83 | 0.65 | 2.69 | 0.1 | 2HWC J1837065 | |
| 3HWC J1843034 | 280.95 | 3.39 | 0.6 | 2.51 | 0.13 | 2HWC J1844032 | |
| 3HWC J1847017 | 282.07 | 1.78 | 0.75 | 2.66 | 0.09 | HESS J1848018 | |
| 3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 2.42 | 0.13 | IGR J184900000 | |
| 3HWC J1857+027 | 284.35 | 2.82 | 0.6 | 2.73 | 0.13 | HESS J1857+026 | |
| 3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 2.44 | 0.09 | MGRO J1908+06 | |
| 3HWC J1912+103 | 288.1 | 10.32 | 0.52 | 2.83 | 0.14 | HESS J1912+101 | |
| 3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 2.69 | 0 | 2HWC J1914+117* | |
| 3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 2.77 | 0 | W 51 | |
| 3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 2.52 | 0.11 | 2HWC J1928+177 | |
| 3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 2.51 | 0.14 | SNR G054.1+00.3 | |
| 3HWC J2019+367 | 304.9 | 36.77 | 0.32 | 2.04 | 0.31 | VER J2019+368 | |
| 3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 2.52 | 0.19 | TeV J2032+4130 | |
| 3HWC J2227+610 | 336.82 | 60.94 | PS | 2.42 | 0 | Boomerang |
4.2 Joint-fit result
We performed the joint fit using the best-fit model of HAWC and added a neutrino source as described in Section 3.4. We found no significant neutrino emission. The smallest pre-trial p-value from the individual source search is 0.07 for 3HWC J2019+367, which corresponds to a post-trial p-value of 0.21 when accounting for the trials in the individual source search. The pre-trial p-value for each source can be found in appendix B. We performed the binomial test and found no significant neutrino emission from any subset of the 22 sources. The pre-trial p-value of the binomial test is 0.09 at k=6 (3HWC J2019+367, 3HWC J2031+415, 3HWC J1912+103, 3HWC J1908+063, 3HWC J1857+027, 3HWC J1928+178) and corresponds to a post-trial p-value of 0.34 when only accounting for trials in the binomial test. Accounting for both the trials from the individual source search and the binomial test, the final post-trial p-value is 0.37.
Based on the null result of the joint search, we also calculated the neutrino flux limits at 90% CL for the 22 sources (see table 3). We account for IceCube systematics by calculating the sensitivity using different Monte Carlo systematic sets and adding the difference to the baseline in quadrature. Appendix C details the effect of IceCube systematics on the flux limit for each source. Table 3 shows the neutrino flux limit for each source. Figure 4 shows, by declination, the neutrino flux limit and the neutrino flux predicted from the best gamma-ray fit assuming all the gamma-ray emission is from hadronic interaction.
| Name | RA | Dec | Extension | Neutrino 90% CL flux limit | Predicted neutrino flux | IceCube 90% energy range |
| [∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV-1cm-2s-1] | [TeV] | |
| 3HWC J0534+220 | 83.64 | 22.01 | PS | 0.3 - 21.6 | ||
| 3HWC J0634+067 | 98.59 | 6.66 | PS | 0.4 - 143.0 | ||
| 3HWC J1757240 | 269.32 | 23.86 | 0.10 | 98.7 - 6870.0 | ||
| 3HWC J1809190 | 272.42 | 19.34 | 0.23 | 38.7 - 468.3 | ||
| 3HWC J1813125 | 273.40 | 12.70 | 0.27 | 41.3 - 3897.3 | ||
| 3HWC J1813174 | 273.40 | 17.74 | 0.26 | 9.60 - 263.5 | ||
| 3HWC J1825134 | 276.49 | 13.63 | 0.46 | 26.2 - 519.7 | ||
| 3HWC J1831095 | 277.85 | 9.85 | 1.14 | 3.1 - 229.4 | ||
| 3HWC J1837066 | 279.38 | 6.83 | 0.65 | 0.7 - 94.1 | ||
| 3HWC J1843034 | 280.95 | 3.39 | 0.60 | 0.7 - 57.0 | ||
| 3HWC J1847017 | 282.07 | 1.78 | 0.75 | 0.5 - 49.1 | ||
| 3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 0.7 - 56.9 | ||
| 3HWC J1857+027 | 284.35 | 2.82 | 0.60 | 0.4 - 29.2 | ||
| 3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 0.6 - 65.1 | ||
| 3HWC J1912+103 | 288.10 | 10.32 | 0.52 | 0.4 - 24.1 | ||
| 3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 0.3 - 73.9 | ||
| 3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 0.3 - 51.8 | ||
| 3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 0.5 - 42.7 | ||
| 3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 0.5 - 36.6 | ||
| 3HWC J2019+367 | 304.90 | 36.77 | 0.32 | 1.1 - 34.2 | ||
| 3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 0.5 - 25.8 | ||
| 3HWC J2227+610 | 336.82 | 60.94 | PS | 0.4 - 78.4 |
Five sources have a lower neutrino 90% CL flux limit than the predicted neutrino flux. In these cases, we conclude that the gamma-ray emission in the TeV range cannot completely originate from hadronic interactions. If we further assume that the leptonic component shares a similar spectral shape in the TeV energy range, we can constrain the percentage of the gamma-ray flux originating from hadronic interactions. Among these sources, 3HWC J2227+610 was studied by the HAWC collaboration and could be associated with supernova remnant G106.3+2.7. HAWC found a hint for PeV proton acceleration and suggested that 3HWC J2227+610 could be a Galactic PeVatron (Albert et al., 2020a). The Crab Nebula (3HWC J0534+220) is the brightest source in the TeV sky and is the most strongly constrained. 3HWC J1922+140 is close to the W 51 region which contains the star-forming region W51B, the supernova remnant W51C, and a potential PWN, CXOU J192318.5+140305 (Abdalla et al., 2018). 3HWC J1922+140 could be a composite of different astrophysical origins. Table 4 shows the constraints on the hadronic fraction, which range from 51% to 85%.
| Name | Neutrino 90% CL flux limit [TeV-1cm-2s-1] | p-value | Hadronic fraction limit | ||
| 3HWC J1847-017 | 2.66 | 0.09 | 1 | 0.79 | |
| 3HWC J1914+118 | 2.69 | 0 | 1 | 0.59 | |
| 3HWC J1922+140 | 2.77 | 0 | 1 | 0.85 | |
| 3HWC J0534+220 | 2.82 | 0.12 | 0.36 | 0.51 | |
| 3HWC J2227+610 | 2.42 | 0 | 1 | 0.82 |
5 Summary and conclusion
In this study, we present a new method for performing multi-messenger spectral fits with the 3ML software framework (Vianello et al., 2017) using HAWC and IceCube data. We performed a joint fit using HAWC and IceCube data to search for neutrino emission from potential Galactic PeVatrons. No significant neutrino emission was observed. The most significant source from the joint fit is 3HWC J2019+367 with the lowest pre-trial p-value of 0.07, corresponding to a post-trial p-value of 0.21. A binomial test for an excess distributed over several sources was performed. The binomial test gave a minimum p-value of 0.09, obtained for the six most significant sources. This corresponds to a post-trial p-value of 0.34. After trial correction for the two searches, the overall post-trial p-value is 0.37. We calculated upper limits at 90% confidence level on the neutrino flux, assuming the spectral form corresponding to the best fit for gamma rays. We found that five sources have a neutrino flux limit that is lower than the predicted neutrino flux of a purely hadronic source. We conclude that these five sources cannot be purely hadronic and constrain the hadronic fraction of these sources. Next-generation neutrino detectors can improve our understanding of the physical processes of these TeV gamma-ray emitters. The proposed IceCube-Gen2 detector (Aartsen et al., 2021), with an increased effective area compared to IceCube, is expected to become up to five times more sensitive to galactic neutrino sources. Several neutrino experiments, such as the proposed Pacific Ocean Neutrino Experiment (P-ONE) (Agostini et al., 2020), Tropical Deep-sea Neutrino Telescope (TRIDENT) (Ye et al., 2023), KM3NeT (currently under construction) (Margiotta, 2014) and Baikal-GVD (operating and undergoing upgrade) (Avrorin et al., 2022), are located in the northern hemisphere and have the potential to detect galactic neutrino sources near the galactic center. The proposed SWGO experiment, which is a ground-based gamma-ray water Cherenkov observatory in the southern hemisphere, would be able to discover more TeV gamma rays sources in the southern sky and candidates of Galactic PeVatrons (Huentemeyer et al., 2019).
Appendix A Likelihood formalism
HAWC performs a binned likelihood analysis for studying gamma-ray sources. In this analysis, we use the energy binning, the fraction of PMTs hit () binning, and also Healpix Nside=1024 spatial binning. We calculate the expected background for each bin by applying direct integration using the zenith distribution of all events passing the gamma/hadron cuts. The technique of direct integration is described in previous HAWC publications (Abeysekara et al., 2019c). The expected excess in every bin is calculated based on the physics model and the detector response of HAWC. The total log-likelihood is the sum of the Poisson log-likelihood of each bin and pixel.
where is the expected background at bin and pixel , is the expected excess per unit flux given the model at bin and pixel , is the flux and is the actual count in bin and pixel . Index loops over the energy and bin and index loops over the number of pixels in the region of interest.
For IceCube, we perform an unbinned likelihood analysis. The total likelihood is given by
where represents the properties of the source hypothesis like location, morphology, and spectrum. is the likelihood of the event . The parameter vector represents the reconstruction information of the event , including reconstructed energy, direction, and estimated angular error. is the total number of events and is the number of signal neutrinos. and are the signal probability density function (PDF) and background probability density function, each consisting of a product of a spatial likelihood and an energy likelihood. A detailed description of the IceCube likelihood for extended sources can be found in Abbasi et al. (2023a). Since the background likelihood only depends on the reconstructed information of the event and does not depend on the source hypothesis, it is a constant during the fitting process. Therefore, we can divide the likelihood by the likelihood of the background-only hypothesis to get the likelihood ratio which is still a valid likelihood. We rewrite the IceCube log-likelihood to be
Since HAWC and IceCube measurements are independent, the total log-likelihood for the joint fit is the addition of the HAWC log-likelihood and IceCube log-likelihood.
The combined log-likelihood is then maximized over the free parameters in the model.
Appendix B Pre-trial p-values of the joint search
Table 5 shows the pre-trial p-value of each source.
| Name | pre-trial p-value | post-trial p-value |
| 3HWC J0534+220 | 0.36 | - |
| 3HWC J0634+067 | 0.19 | - |
| 3HWC J1757-240 | 1 | - |
| 3HWC J1809-190 | 1 | - |
| 3HWC J1813-125 | 1 | - |
| 3HWC J1813-174 | 0.70 | - |
| 3HWC J1825-134 | 1 | - |
| 3HWC J1831-095 | 1 | - |
| 3HWC J1837-066 | 1 | - |
| 3HWC J1843-034 | 1 | - |
| 3HWC J1847-017 | 1 | - |
| 3HWC J1849+001 | 1 | - |
| 3HWC J1857+027 | 0.13 | - |
| 3HWC J1908+063 | 0.10 | - |
| 3HWC J1912+103 | 0.08 | - |
| 3HWC J1914+118 | 1 | - |
| 3HWC J1922+140 | 1 | - |
| 3HWC J1928+178 | 0.14 | - |
| 3HWC J1930+188 | 1 | - |
| 3HWC J2019+367 | 0.07 | 0.21 |
| 3HWC J2031+415 | 0.08 | - |
| 3HWC J2227+610 | 1 | - |
Appendix C IceCube Systematics
In this analysis, we consider the effect of uncertainties on the scattering and absorption of light in the ice, uncertainty in DOM efficiency, and uncertainty in the modeling of ice properties. We consider uncertainty in absorption and scattering, in DOM efficiency, and uncertainty in hole ice modeling parameters as described in Eller et al. (2023). The method of handling the systematic uncertainties is similar to the one previously used in IceCube’s measurement of the diffuse neutrino flux (Abbasi et al., 2022b). We use different simulation data sets, each with the value of one parameter changed to account for its systematic uncertainty. The change in sensitivity for the joint analysis as compared to the baseline was calculated for the different sources of systematic uncertainties, and these changes were added in quadrature. Table 6 shows the total systematic uncertainty on the neutrino 90% CL flux limit from each source. The final neutrino flux limit of each source is increased by the total systematic uncertainty and is shown in table 3.
| Name | Total systematic uncertainties |
| 3HWC J0534+220 | 18% |
| 3HWC J0634+067 | 25% |
| 3HWC J1757-240 | 25% |
| 3HWC J1809-190 | 39% |
| 3HWC J1813-125 | 27% |
| 3HWC J1813-174 | 54% |
| 3HWC J1825-134 | 30% |
| 3HWC J1831-095 | 40% |
| 3HWC J1837-066 | 33% |
| 3HWC J1843-034 | 19% |
| 3HWC J1847-017 | 21% |
| 3HWC J1849+001 | 19% |
| 3HWC J1857+027 | 18% |
| 3HWC J1908+063 | 18% |
| 3HWC J1912+103 | 21% |
| 3HWC J1914+118 | 20% |
| 3HWC J1922+140 | 14% |
| 3HWC J1928+178 | 13% |
| 3HWC J1930+188 | 23% |
| 3HWC J2019+367 | 14% |
| 3HWC J2031+415 | 16% |
| 3HWC J2227+610 | 30% |

### v2

@@
-Search for joint multimessenger signals from potential Galactic PeVatrons with HAWC and IceCube
+Search for joint multimessenger signals from potential galactic cosmic-ray accelerators with HAWC and IceCube
 Abstract
-Galactic PeVatrons are sources that can accelerate cosmic rays to PeV energies.
-The high-energy cosmic rays are expected to interact with the surrounding ambient material or radiation, resulting in the production of gamma rays and neutrinos.
+The origin of high-energy galactic cosmic rays is yet to be understood, but some galactic cosmic-ray accelerators can accelerate cosmic rays up to PeV energies.
+The high-energy cosmic rays are expected to interact with the surrounding material or radiation, resulting in the production of gamma rays and neutrinos.
 To optimize for the detection of such associated production of gamma rays and neutrinos for a given source morphology and spectrum, a multi-messenger analysis that combines gamma rays and neutrinos is required.
@@
 We report the best-fit gamma-ray model and 90% CL neutrino flux limit from the 22 sources.
-From the neutrino flux limit, we conclude that the gamma-ray emission from five of the sources can not be produced purely from hadronic interactions.
+From the neutrino flux limit, we conclude that for five of the sources, the gamma-ray emission observed by HAWC cannot be produced purely from hadronic interactions.
 We report the limit for the fraction of gamma rays produced by hadronic interactions for these five sources.
 1 Introduction
-Since the discovery of cosmic rays (CRs) more than 100 years ago (Hess, 1912), but where these high-energy particles come from has not yet been well understood.
-The highest energy charged particles, protons and nuclei, are thought to be extraterrestrial in origin, and their spectrum can be parametrized using a power law with an index of 2.7 up to PeV (or the “knee” in the spectrum), above which the spectrum steepens (Abbasi et al., 2018).
+Cosmic rays (CRs) were discovered more than 100 years ago (Hess, 1912), but where these high-energy particles come from has not yet been well understood.
+The highest energy charged particles, protons and nuclei, are thought to be extraterrestrial in origin, and their spectrum can be parametrized using a power law with an index of 2.7 up to PeV (or the “knee” in the spectrum), above which the spectrum steepens (Nagano et al., 1984; Abbasi et al., 2018; Cao et al., 2024a; Amenomori et al., 2008; Aartsen et al., 2020a; Glasmacher et al., 1999).
 The CRs below the knee are thought to originate from sources within the Milky Way galaxy, hinting at the presence of PeV hadron accelerators within the galaxy known as Galactic PeVatrons (Gabici & Aharonian, 2007).
-Galactic PeVatrons are expected to accelerate particles up to PeV energies through Fermi acceleration and these accelerated particles are expected to be confined within the Milky Way galaxy (Abbasi et al., 2018).
+Galactic PeVatrons are expected to accelerate particles up to PeV energies through Fermi acceleration and these accelerated particles are expected to be confined within the Milky Way galaxy (de Oña Wilhelmi et al., 2024).
+The AS gamma collaboration reported the detection of diffuse gamma rays from the galactic disk with energy between 100 TeV and 1 PeV, hinting at the presence of PeVatrons within our galaxy (Amenomori et al., 2021).
 Supernova remnants (SNRs) are a potential class of Galactic PeVatrons.
 The particles are accelerated by diffusive shock acceleration in the expanding shock.
-Only 10% of the energy of SNRs is needed to explain the observed CR energy density in the Galaxy (Blandford & Ostriker, 1978).
+Only 10% of the energy of SNRs is needed to account for the observed CR energy density in the Galaxy (Blandford & Ostriker, 1978).
 For the most energetic galactic CRs, it is assumed that only young SNRs can accelerate CRs beyond 100 TeV (Bell et al., 2013).
@@
 High-energy gamma rays and neutrinos can be efficiently produced by the decay of charged and neutral pions.
-Other potential source classes of Galactic PeVatrons include young massive star clusters, YMCs (Bykov et al., 2020) and pulsar wind nebulae, PWNe (Di Palma et al., 2017).
+Other potential classes of Galactic PeVatrons include young massive star clusters, YMCs (Bykov et al., 2020) and pulsar wind nebulae, PWNe (Di Palma et al., 2017).
 TeV gamma rays can also be produced by leptonic interactions such as inverse Compton scattering.
@@
 The HAWC collaboration also reported HAWC J2227+610 as a potential PeVatron candidate from TeV gamma-ray observations and concluded the proton cutoff energy is at least 800 TeV (Albert et al., 2020a).
-The HAWC collaboration studied the ultra-high-energy spectrum off MGRO J1908+06 and their model allows a hadronic component at the highest energy (Albert et al., 2022).
+The HAWC collaboration studied the ultra-high-energy spectrum of MGRO J1908+06 and their model allows a hadronic component at the highest energy (Albert et al., 2022).
 The HESS Collaboration reported evidence of a PeVatron at the galactic center based on gamma-ray observations up to 40 TeV (Abramowski et al., 2016).
+The LHAASO collaboration reported an ultrahigh-energy gamma ray bubble up to a few PeV from the direction of the Cygnus X region, and suggested that a super PeVatron with maximum proton energy of 10 PeV could power the source (LHAASO Collaboration, 2024).
+The LHAASO collaboration detected 12 gamma-ray sources with emissions above 100 TeV (Cao et al., 2021) and recently reported their first catalog 1LHAASO (Cao et al., 2024b).
+The 1LHAASO contains 90 gamma-ray sources, and 43 sources have emission above 100 TeV, making them possible Galactic PeVatrons.
+Many efforts have been made to search for PeVatrons using neutrino data.
+Huang & Li (2022) set a constraint on neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using previous IceCube neutrino upper limits.
+They used measurements from other gamma-ray instruments and constrained the neutrino flux using the measured gamma-ray spectrum.
+Another search by Huang & Li (2022) used 10 years of public IceCube track data to search for neutrino emission from the same LHAASO catalog (Cao et al., 2021) with fixed morphology and different spectral models.
 The IceCube Neutrino Observatory has previously searched for neutrino emission from galactic sources, both individual sources and classes of sources.
 Abbasi et al.
-(2023a) searched for neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using a fixed spatial extension and a power law spectrum.
+(2023a) searched for neutrino emission from LHAASO’s first catalog of sources with energy above 100 TeV (Cao et al., 2021) using a fixed spatial extension and a power law spectrum and 12 years of IceCube track data.
 They found no significant emission, neither from single sources nor from stacked catalogs of SNRs and PWNe.
@@
 However, these searches have not directly incorporated gamma-ray data.
-The simultaneous use of both data sets incorporates more physically motivated spectra and provides a direct constraint on the hadronic ratio, i.e.
-the fraction of gamma rays originating from hadronic interactions.
+The simultaneous use of both data sets incorporates more physically motivated spectra and more accurate morphology.
+In this work, for the first time, we simultaneously use both gamma-ray and neutrino data to improve the sensitivity of the search and provide a direct constraint on the hadronic ratio, i.e., the fraction of gamma rays originating from hadronic interactions.
+The improvement in sensitivity of the search comes from the additional 2 years for data taken with full detector compared to Abbasi et al.
+(2023a) and the modeling of the neutrino emission under the assumption that the gamma rays and neutrinos originated from the same population of hadrons.
 If the gamma rays and neutrinos originate from hadronic interactions, the spectra of the gamma rays and neutrinos have a fixed relationship.
@@
 The events within an analysis bin are expected to have similar characteristics, including angular resolution, primary gamma-ray energy, and gamma-to-hadron likelihood (signal-to-noise ratio), which allows HAWC to perform a binned likelihood analysis.
-In this study, we use 2141 days of HAWC data using the Pass 5 reconstruction (Yun-Carcamo et al., 2023).
+In this study, we use 2141 days of HAWC data using the Pass 5 reconstruction (Albert et al., 2024).
 The event-related energy is estimated using the “ground parameter” estimator described in Abeysekara et al.
@@
 The dataset covers the period from June 15, 2015 to October 21, 2021 with a livetime of more than 90%.
+The dataset contains 29% more data with an improved reconstruction compared to the dataset used in the third HAWC catalog (3HWC) (Albert et al., 2020b).
 2.2 The IceCube Neutrino Observatory
@@
 3.1 Source selection
-HAWC detected 65 sources with a significance above in the third HAWC catalog (3HWC) (Albert et al., 2020b) using 1523 days of data.
+The third HAWC catalog (3HWC) includes 65 sources with a significance above 5 based on 1523 days of data (Albert et al., 2020b).
 The 3HWC provides the location and the gamma-ray flux at a pivot energy of 7 TeV for each source in the catalog.
@@
 Signal trials are generated by scrambling data and adding signal using simulation data.
-The calculation of sensitivity is based on a maximum likelihood method as used in the IceCube 10-year search for neutrino sources(Aartsen et al., 2020a).
+The calculation of sensitivity is based on a maximum likelihood method as used in the IceCube 10-year search for neutrino sources(Aartsen et al., 2020b).
 In this work, we use a new likelihood analysis software component called i3mla (Abbasi et al., 2021) with the multi-mission maximum likelihood framework (3ML) (Vianello et al., 2017).
@@
 We select sources whose predicted neutrino flux from hadronic interactions is larger than the sensitivity of the IceCube point source search.
-The hard spectrum is an optimistic choice made to avoid missing any potential sources in our source selection step.
-The final sensitivity of the joint search will change depending on the best-fit gamma-ray spectrum and spatial extension.
+The source selection step is motivated by the trade-off between searching as many possible sources as possible to avoid missing potential sources and avoiding suffering sensitivity loss from the trial factor.
+The hard spectrum is an optimistic choice made to avoid missing any potential sources in our source selection step since it is close to the hardest spectrum expected.
+The sensitivity of the joint search will depend on the best-fit gamma-ray spectrum and spatial extension and is different from the point source sensitivity here.
 The sensitivity of IceCube and the predicted neutrino flux from 3HWC sources are shown in Figure 1.
@@
 We use 2141 days of HAWC ground parameter energy estimator maps with the Pass 5 reconstruction for our model selection procedure.
-The dataset contains more data and a better energy estimator compared to the dataset used in 3HWC.
+The dataset contains 29% more data with an improved reconstruction compared to the dataset used in 3HWC.
 The new reconstruction provides a better angular reconstruction and gamma/hadron separation.
@@
 The method is conceptually identical to the previous IceCube 10-year time-integrated point source search in Aartsen et al.
-(2020a) and the extended source search in Abbasi et al.
+(2020b) and the extended source search in Abbasi et al.
 (2023a, c).
@@
 Flux normalization | Float | Float | Independent |
-To calculate the p-value, we perform a large number of background trials by scrambling the IceCube data in RA and building the background test statistic (TS) distribution (Aartsen et al., 2020b).
+To calculate the p-value, we perform a large number of background trials by scrambling the IceCube data in RA and building the background test statistic (TS) distribution (Aartsen et al., 2020c).
 This is done to avoid any dependence on the simulation when testing for a neutrino source.
@@
 We have developed a new custom plugin for 3ML to read the likelihood table and construct a spline.
-The 3ML plugin returns the spline value for the likelihood evaluation.
+The 3ML plugin then returns the spline values for the likelihood evaluations.
 This method speeds up the likelihood calculation because the program does not have to calculate the HAWC likelihood for each pixel and each bin during the fitting.
@@
 For each background trial, the IceCube data are scrambled in RA while keeping the HAWC component of the likelihood unchanged, and the joint analysis is performed.
-To determine the sensitivity to the neutrino flux of the joint search, we inject different simulated values of the neutrino flux at the location of the best-fit gamma-ray source with its best-fit extension and spectral shape.
+To determine the sensitivity of the joint search for a source with the best-fit parameters, we inject different simulated values of the neutrino flux at the location of the best-fit gamma-ray source with its best-fit extension and spectral shape.
 The 90% CL sensitivity is the neutrino flux required for the TS to exceed the median of the background-only TS distribution 90% of the time.
@@
 The errors of each type of systematic uncertainty are added in quadrature to obtain the total systematic uncertainty, which is given in the table 2.
-| Name | RA [∘] | Dec [∘] | Ext [∘] | [TeV-1cm-2s-1] | Nearest TeVCat source | ||
-| 3HWC J0534+220 | 83.64 | 22.01 | PS | 2.82 | 0.12 | Crab | |
-| 3HWC J0634+067 | 98.59 | 6.66 | 0.49 | 2.57 | 0 | HAWC J0635+070 | |
-| 3HWC J1757240 | 269.32 | 23.86 | PS | 2.58 | 0 | HESS J1800240B | |
-| 3HWC J1809190 | 272.42 | 19.34 | 0.23 | 2.09 | 0.23 | HESS J1809193 | |
-| 3HWC J1813125 | 273.4 | 12.7 | 0.27 | 2.6 | 0 | HESS J1813-126 | |
-| 3HWC J1813174 | 273.4 | 17.74 | 0.26 | 1.91 | 0.35 | 2HWC J1814173 | |
-| 3HWC J1825134 | 276.49 | 13.63 | 0.46 | 2.45 | 0.13 | 2HWC J1825134 | |
-| 3HWC J1831095 | 277.85 | 9.85 | 1.14 | 2.59 | 0.14 | HESS J1831098 | |
-| 3HWC J1837066 | 279.38 | 6.83 | 0.65 | 2.69 | 0.1 | 2HWC J1837065 | |
-| 3HWC J1843034 | 280.95 | 3.39 | 0.6 | 2.51 | 0.13 | 2HWC J1844032 | |
-| 3HWC J1847017 | 282.07 | 1.78 | 0.75 | 2.66 | 0.09 | HESS J1848018 | |
-| 3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 2.42 | 0.13 | IGR J184900000 | |
-| 3HWC J1857+027 | 284.35 | 2.82 | 0.6 | 2.73 | 0.13 | HESS J1857+026 | |
-| 3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 2.44 | 0.09 | MGRO J1908+06 | |
-| 3HWC J1912+103 | 288.1 | 10.32 | 0.52 | 2.83 | 0.14 | HESS J1912+101 | |
-| 3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 2.69 | 0 | 2HWC J1914+117* | |
-| 3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 2.77 | 0 | W 51 | |
-| 3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 2.52 | 0.11 | 2HWC J1928+177 | |
-| 3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 2.51 | 0.14 | SNR G054.1+00.3 | |
-| 3HWC J2019+367 | 304.9 | 36.77 | 0.32 | 2.04 | 0.31 | VER J2019+368 | |
-| 3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 2.52 | 0.19 | TeV J2032+4130 | |
-| 3HWC J2227+610 | 336.82 | 60.94 | PS | 2.42 | 0 | Boomerang |
+We also reported the energy range of each source.
+The upper end of this range is defined such that if the spectrum is cut off above this energy the significance of the source drops by .
+The corresponding definition is used for the lower end of the sensitivity range.
+The threshold is chosen to reasonably claim that HAWC data can probe the quoted energy range at a significance of or at the 95% confidence level.
+Name | RA | Dec | Ext | Nearest TeVCat | Energy range | |||
+[∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV] | ||||
+3HWC J0534+220 | 83.64 | 22.01 | PS | 2.82 | 0.12 | Crab | 0.1 - 106.1 | |
+3HWC J0634+067 | 98.59 | 6.66 | 0.49 | 2.57 | 0 | HAWC J0635+070 | 1.7-53.1 | |
+3HWC J1757240 | 269.32 | 23.86 | PS | 2.58 | 0 | HESS J1800240B | 16.9 - 59.7 | |
+3HWC J1809190 | 272.42 | 19.34 | 0.23 | 2.09 | 0.23 | HESS J1809193 | 3.0 - 94.5 | |
+3HWC J1813125 | 273.4 | 12.7 | 0.27 | 2.6 | 0 | HESS J1813-126 | 2.4 - 59.6 | |
+3HWC J1813174 | 273.4 | 17.74 | 0.26 | 1.91 | 0.35 | 2HWC J1814173 | 3.4 - 106.0 | |
+3HWC J1825134 | 276.49 | 13.63 | 0.46 | 2.45 | 0.13 | 2HWC J1825134 | 0.5 - 106.0 | |
+3HWC J1831095 | 277.85 | 9.85 | 1.14 | 2.59 | 0.14 | HESS J1831098 | 0.4 - 118.7 | |
+3HWC J1837066 | 279.38 | 6.83 | 0.65 | 2.69 | 0.1 | 2HWC J1837065 | 0.2 - 133.2 | |
+3HWC J1843034 | 280.95 | 3.39 | 0.6 | 2.51 | 0.13 | 2HWC J1844032 | 0.3 - 75.0 | |
+3HWC J1847017 | 282.07 | 1.78 | 0.75 | 2.66 | 0.09 | HESS J1848018 | 0.3 - 84.0 | |
+3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 2.42 | 0.13 | IGR J184900000 | 0.4 - 149.8 | |
+3HWC J1857+027 | 284.35 | 2.82 | 0.6 | 2.73 | 0.13 | HESS J1857+026 | 0.3 - 133.2 | |
+3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 2.44 | 0.09 | MGRO J1908+06 | 0.4 - 168.0 | |
+3HWC J1912+103 | 288.1 | 10.32 | 0.52 | 2.83 | 0.14 | HESS J1912+101 | 0.4 - 94.4 | |
+3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 2.69 | 0 | 2HWC J1914+117* | 0.3 - 106.0 | |
+3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 2.77 | 0 | W 51 | 0.2 - 47.3 | |
+3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 2.52 | 0.11 | 2HWC J1928+177 | 0.6 - 106.0 | |
+3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 2.51 | 0.14 | SNR G054.1+00.3 | 0.7 - 105.9 | |
+3HWC J2019+367 | 304.9 | 36.77 | 0.32 | 2.04 | 0.31 | VER J2019+368 | 0.9 - 106.0 | |
+3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 2.52 | 0.19 | TeV J2032+4130 | 0.5 - 105.9 | |
+3HWC J2227+610 | 336.82 | 60.94 | PS | 2.42 | 0 | Boomerang | 16.7 - 33.5 |
 4.2 Joint-fit result
@@
 The smallest pre-trial p-value from the individual source search is 0.07 for 3HWC J2019+367, which corresponds to a post-trial p-value of 0.21 when accounting for the trials in the individual source search.
-The pre-trial p-value for each source can be found in appendix B.
 We performed the binomial test and found no significant neutrino emission from any subset of the 22 sources.
-The pre-trial p-value of the binomial test is 0.09 at k=6 (3HWC J2019+367, 3HWC J2031+415, 3HWC J1912+103, 3HWC J1908+063, 3HWC J1857+027, 3HWC J1928+178) and corresponds to a post-trial p-value of 0.34 when only accounting for trials in the binomial test.
+The smallest pre-trial p-value of the binomial test is 0.09 at =6 (3HWC J2019+367, 3HWC J2031+415, 3HWC J1912+103, 3HWC J1908+063, 3HWC J1857+027, 3HWC J1928+178) and corresponds to a post-trial p-value of 0.34 when only accounting for trials in the binomial test.
 Accounting for both the trials from the individual source search and the binomial test, the final post-trial p-value is 0.37.
+We also computed the 90% neutrino energy range for each source using the best-fit morphology and spectrum.
+The 90% neutrino energy range is the central 90% range of true neutrino energy when injecting signal neutrinos with the source’s best-fit gamma-ray morphology and spectrum.
+IceCube applies a higher energy threshold for neutrino events from the southern sky due to the containment of atmospheric muon, resulting in a higher energy range for southern sources.
 Based on the null result of the joint search, we also calculated the neutrino flux limits at 90% CL for the 22 sources (see table 3).
 We account for IceCube systematics by calculating the sensitivity using different Monte Carlo systematic sets and adding the difference to the baseline in quadrature.
-Appendix C details the effect of IceCube systematics on the flux limit for each source.
+Appendix B details the effect of IceCube systematics on the flux limit for each source.
 Table 3 shows the neutrino flux limit for each source.
 Figure 4 shows, by declination, the neutrino flux limit and the neutrino flux predicted from the best gamma-ray fit assuming all the gamma-ray emission is from hadronic interaction.
-| Name | RA | Dec | Extension | Neutrino 90% CL flux limit | Predicted neutrino flux | IceCube 90% energy range |
-| [∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV-1cm-2s-1] | [TeV] | |
-| 3HWC J0534+220 | 83.64 | 22.01 | PS | 0.3 - 21.6 | ||
-| 3HWC J0634+067 | 98.59 | 6.66 | PS | 0.4 - 143.0 | ||
-| 3HWC J1757240 | 269.32 | 23.86 | 0.10 | 98.7 - 6870.0 | ||
-| 3HWC J1809190 | 272.42 | 19.34 | 0.23 | 38.7 - 468.3 | ||
-| 3HWC J1813125 | 273.40 | 12.70 | 0.27 | 41.3 - 3897.3 | ||
-| 3HWC J1813174 | 273.40 | 17.74 | 0.26 | 9.60 - 263.5 | ||
-| 3HWC J1825134 | 276.49 | 13.63 | 0.46 | 26.2 - 519.7 | ||
-| 3HWC J1831095 | 277.85 | 9.85 | 1.14 | 3.1 - 229.4 | ||
-| 3HWC J1837066 | 279.38 | 6.83 | 0.65 | 0.7 - 94.1 | ||
-| 3HWC J1843034 | 280.95 | 3.39 | 0.60 | 0.7 - 57.0 | ||
-| 3HWC J1847017 | 282.07 | 1.78 | 0.75 | 0.5 - 49.1 | ||
-| 3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 0.7 - 56.9 | ||
-| 3HWC J1857+027 | 284.35 | 2.82 | 0.60 | 0.4 - 29.2 | ||
-| 3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 0.6 - 65.1 | ||
-| 3HWC J1912+103 | 288.10 | 10.32 | 0.52 | 0.4 - 24.1 | ||
-| 3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 0.3 - 73.9 | ||
-| 3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 0.3 - 51.8 | ||
-| 3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 0.5 - 42.7 | ||
-| 3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 0.5 - 36.6 | ||
-| 3HWC J2019+367 | 304.90 | 36.77 | 0.32 | 1.1 - 34.2 | ||
-| 3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 0.5 - 25.8 | ||
-| 3HWC J2227+610 | 336.82 | 60.94 | PS | 0.4 - 78.4 |
+Name | RA | Dec | Extension | p-value | Neutrino 90% CL flux limit | Predicted neutrino flux | Energy range |
+[∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV-1cm-2s-1] | [TeV] | ||
+3HWC J0534+220 | 83.64 | 22.01 | PS | 0.36 | 0.3 - 21.6 | ||
+3HWC J0634+067 | 98.59 | 6.66 | PS | 0.19 | 0.4 - 143.0 | ||
+3HWC J1757240 | 269.32 | 23.86 | 0.10 | 1 | 98.7 - 6870.0 | ||
+3HWC J1809190 | 272.42 | 19.34 | 0.23 | 1 | 38.7 - 468.3 | ||
+3HWC J1813125 | 273.40 | 12.70 | 0.27 | 1 | 41.3 - 3897.3 | ||
+3HWC J1813174 | 273.40 | 17.74 | 0.26 | 0.7 | 9.60 - 263.5 | ||
+3HWC J1825134 | 276.49 | 13.63 | 0.46 | 1 | 26.2 - 519.7 | ||
+3HWC J1831095 | 277.85 | 9.85 | 1.14 | 1 | 3.1 - 229.4 | ||
+3HWC J1837066 | 279.38 | 6.83 | 0.65 | 1 | 0.7 - 94.1 | ||
+3HWC J1843034 | 280.95 | 3.39 | 0.60 | 1 | 0.7 - 57.0 | ||
+3HWC J1847017 | 282.07 | 1.78 | 0.75 | 1 | 0.5 - 49.1 | ||
+3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 1 | 0.7 - 56.9 | ||
+3HWC J1857+027 | 284.35 | 2.82 | 0.60 | 0.13 | 0.4 - 29.2 | ||
+3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 0.1 | 0.6 - 65.1 | ||
+3HWC J1912+103 | 288.10 | 10.32 | 0.52 | 0.08 | 0.4 - 24.1 | ||
+3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 1 | 0.3 - 73.9 | ||
+3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 1 | 0.3 - 51.8 | ||
+3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 0.14 | 0.5 - 42.7 | ||
+3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 0.1 | 0.5 - 36.6 | ||
+3HWC J2019+367 | 304.90 | 36.77 | 0.32 | 0.07 | 1.1 - 34.2 | ||
+3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 0.08 | 0.5 - 25.8 | ||
+3HWC J2227+610 | 336.82 | 60.94 | PS | 1 | 0.4 - 78.4 |
 Five sources have a lower neutrino 90% CL flux limit than the predicted neutrino flux.
-In these cases, we conclude that the gamma-ray emission in the TeV range cannot completely originate from hadronic interactions.
+In these cases, we conclude that the gamma-ray emission detected in the HAWC energy range cannot completely originate from hadronic interactions.
+This assumes that the HAWC observed spectrum extends to the IceCube energy range and it doesn’t account for HAWC spectral uncertainty.
 If we further assume that the leptonic component shares a similar spectral shape in the TeV energy range, we can constrain the percentage of the gamma-ray flux originating from hadronic interactions.
@@
 The Crab Nebula (3HWC J0534+220) is the brightest source in the TeV sky and is the most strongly constrained.
+However, the Crab Nebula is widely believed to be a leptonic source.
+A more sensitive neutrino instrument is required to put a relevant constraint, in the 1% range, on the hadronic fraction.
 3HWC J1922+140 is close to the W 51 region which contains the star-forming region W51B, the supernova remnant W51C, and a potential PWN, CXOU J192318.5+140305 (Abdalla et al., 2018).
@@
 3HWC J2227+610 | 2.42 | 0 | 1 | 0.82 |
-5 Summary and conclusion
-In this study, we present a new method for performing multi-messenger spectral fits with the 3ML software framework (Vianello et al., 2017) using HAWC and IceCube data.
-We performed a joint fit using HAWC and IceCube data to search for neutrino emission from potential Galactic PeVatrons.
+5 Summary and discussion
+In this paper, we present a new method for performing multi-messenger spectral fits with the 3ML software framework (Vianello et al., 2017) using HAWC and IceCube data.
+This is motivated by the predicted neutrino emission by Galactic cosmic-ray accelerators with gamma-ray emission.
+The new joint-fitting method with gamma-ray data and neutrino data brings improvement in search sensitivity due to more detailed modeling of the neutrino emission under the assumption that the gamma rays and neutrinos originated from the same population of hadrons.
+In addition, we could provide a direct constraint on the hadronic ratio in the case of null detection using the emission model derived from the gamma-ray data if the assumption is true.
+We stress the caveat that a mis-modeling of the emission region, for example, due to gamma rays and neutrinos being produced in different regions of a source, could weaken the constraints presented in this paper.
+We used the HAWC data to obtain the best-fit model for the gamma-ray emission.
+We performed a joint fit using HAWC and IceCube data to search for neutrino emission from potential Galactic cosmic-ray accelerators.
 No significant neutrino emission was observed.
@@
 A binomial test for an excess distributed over several sources was performed.
-The binomial test gave a minimum p-value of 0.09, obtained for the six most significant sources.
-This corresponds to a post-trial p-value of 0.34.
+The binomial test gave a pre-trial p-value of 0.09, obtained for the six most significant sources and corresponds to a post-trial p-value of 0.34.
 After trial correction for the two searches, the overall post-trial p-value is 0.37.
@@
 We conclude that these five sources cannot be purely hadronic and constrain the hadronic fraction of these sources.
+We found that for some of the sources, the HAWC data do not show a 2 significance for emission above 100 TeV, meaning that the analysis would not warrant the conclusion that those sources are PeVatrons even if neutrinos are detected from the source.
+In addition, the neutrino 90% energy range for some of the sources does not exceed 50 TeV, meaning the neutrino data would likely not be grounds to establish PeV cosmic-ray acceleration in the case of detection.
+A more sensitive experiment such as LHAASO could potentially determine the limit on the maximum energy of the source, and neutrino detection could act as evidence of hadronic interactions.
+In addition, it could provide more PeVatron candidates with better gamma-ray spectrum and morphology measurement.
+The recently released 1LHAASO catalog contains 43 sources with gamma-ray emission above 100 TeV (Cao et al., 2024b) and a joint search with neutrino data will provide better sensitivity and constraints on their hadronic emission.
 Next-generation neutrino detectors can improve our understanding of the physical processes of these TeV gamma-ray emitters.
@@
 The combined log-likelihood is then maximized over the free parameters in the model.
-Appendix B Pre-trial p-values of the joint search
-Table 5 shows the pre-trial p-value of each source.
-| Name | pre-trial p-value | post-trial p-value |
-| 3HWC J0534+220 | 0.36 | - |
-| 3HWC J0634+067 | 0.19 | - |
-| 3HWC J1757-240 | 1 | - |
-| 3HWC J1809-190 | 1 | - |
-| 3HWC J1813-125 | 1 | - |
-| 3HWC J1813-174 | 0.70 | - |
-| 3HWC J1825-134 | 1 | - |
-| 3HWC J1831-095 | 1 | - |
-| 3HWC J1837-066 | 1 | - |
-| 3HWC J1843-034 | 1 | - |
-| 3HWC J1847-017 | 1 | - |
-| 3HWC J1849+001 | 1 | - |
-| 3HWC J1857+027 | 0.13 | - |
-| 3HWC J1908+063 | 0.10 | - |
-| 3HWC J1912+103 | 0.08 | - |
-| 3HWC J1914+118 | 1 | - |
-| 3HWC J1922+140 | 1 | - |
-| 3HWC J1928+178 | 0.14 | - |
-| 3HWC J1930+188 | 1 | - |
-| 3HWC J2019+367 | 0.07 | 0.21 |
-| 3HWC J2031+415 | 0.08 | - |
-| 3HWC J2227+610 | 1 | - |
-Appendix C IceCube Systematics
+Appendix B IceCube Systematics
 In this analysis, we consider the effect of uncertainties on the scattering and absorption of light in the ice, uncertainty in DOM efficiency, and uncertainty in the modeling of ice properties.